# Amplifying hybrid entangled states and superpositions of coherent states

### v1

@@ V Conclusion
We have investigated two amplification schemes, the photon-addition-and-then-subtraction scheme and the $(\hat{a}^{\dagger})^{2}$ scheme, applied to two types of nonclassical states, i.e., hybrid states and SCSs. We have obtained fidelity, gain, and quantum Fisher information of each scheme for each state. For both the schemes, the photon addition and then subtraction scheme has the advantage of higher fidelity, thus it is appropriate for amplifying states with relatively low amplitudes. On the contrary, the $(\hat{a}^{\dagger})^{2}$ scheme has the advantage of higher gain, and it is better for amplifying states with intermediate amplitudes. For sufficiently large amplitudes, the difference between the two schemes vanishes.
@@
For hybrid states, the fidelity and gain under both the schemes are the same as those for the case of coherent states, independent of dimension $d$ or qudit number $k$ due to the orthogonality of the DV part {|$\{\ket{n}\}_{n}$. We have obtained quantum Fisher information of hybrid states that reflects the overall performance of the noiseless amplification. Our analysis shows that the photon addition and then subtraction scheme is better for amplitude between 00 and $0.9$, and the $(\hat{a}^{\dagger})^{2}$ scheme is better for amplitude larger than $0.9$.
@@
For SCSs, behaviors of the fidelity and gain depend on dimension $d$ and qudit number $k$. In a fidelity graph, a smaller qudit number $k$ shows trough(s) for smaller amplitude with larger depth. In gain graphs, similar behavior is observed but two states with the last two qudit numbers $d-2$ and $d-1$ act very differently. We have analyzed the graph focusing on the pseudo-number property of the superposition of coherent states. Unlike hybrid states, quantum Fisher information graphs show fluctuations as amplitude increases and thus do not reflect the overall performance of noiseless amplification except $d=2$. Consequently, quantum Fisher information is adopted as an example of the performance of amplified superposition of coherent states, and consistent criterion extended from even and odd cat states, to compare with other optical states. Since fidelity and gain graphs are not trivial, and quantum Fisher information does not indicate overall performance, an appropriate amplification scheme should be chosen considering the goals and conditions of initial states.
@@ Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (Nos. RS-2024-00413957, RS-2024-00438415, NRF-2023R1A2C1006115, RS-2024-00437191, and NRF-2022M3K4A1097117) and by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2021-0-01059 and IITP-2023-2020-0-01606).

### v2

@@ VI Conclusions
We have obtained the fidelity, the optimized gain, and the quantum Fisher information of each scheme for each state. For both states, the $\hat{a}\hat{a}^{\dagger}{}$ scheme has the advantage of higher fidelity, and it is appropriate for amplifying states with relatively small amplitudes ($\alpha\lesssim 1.5$). On the contrary, the $\hat{a}^{\dagger}{}^{2}$ scheme has the advantage of higher gain, and it is better for amplifying states with intermediate amplitudes ($1.5\lesssim\alpha\lesssim 2.5$). For sufficiently large amplitudes, the difference between the two schemes vanishes.
@@
For the HESs, the fidelity and gain under both schemes are the same as those for the case of coherent states, independent of dimension $d$ or qudit number $k$ due to the orthogonality of the DV part. We have obtained the quantum Fisher information of the HESs that reflects the overall performance of the noiseless amplification. Our analysis shows that the $\hat{a}^{\dagger}\hat{a}$ scheme is better for $\alpha\lesssim 0.9$, while the $\hat{a}^{\dagger}{}^{2}$ scheme works generally better for $\alpha$ larger than $0.9$.
@@
For the SCSs, the behaviors of the fidelity and gain for the amplified states depend on dimension $d$ and qudit number $k$. In the fidelity graphs, a smaller qudit number $k$ shows trough(s) for a smaller amplitude with larger depth. In the gain graphs, similar behaviors are observed but two states with the last two qudit numbers $d-2$ and $d-1$ act very differently. We have analyzed the graphs focusing on the pseudo-number property of the superposition of coherent states. Unlike the HESs, the quantum Fisher information graphs show fluctuations as the amplitude increases and thus do not reflect the overall performance of noiseless amplification except $d=2$. Consequently, the quantum Fisher information is adopted as an example of the performance of the amplified SCS, and consistent criterion extended from even and odd cat states, to compare with other optical states. Since the fidelity and gain graphs are not trivial, and the quantum Fisher information does not indicate overall performance, an appropriate amplification scheme should be chosen considering the goals and conditions of initial states.
@@ Acknowledgements
The first two authors contributed equally to this work. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (Nos. RS-2024-00413957, RS-2024-00438415, NRF-2023R1A2C1006115, and RS-2024-00437191) and by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2024-2020-0-01606). This work was also supported by the Institute of Applied Physics, Seoul National University.